A casino offers the following game: you draw one card from a standard 52-card deck. If you draw a jack, you win $1.25. If you draw a queen, you win $3.25. If you draw a king, you win $5.25 dollars. If you draw any ace except the ace of spades, you win $6.75. If you draw the ace of spades, you win $8.5.

The entry fee to play this game is $2. Compute the expected value of this gamble (include the entry fee in your expected value).
1
Expert's answer
2020-12-08T06:22:33-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x& p(x) \\ \hline
 -2 & {9 \over 13} \\
 -0.75 & {1 \over 13} \\
1.25 & {1 \over 13} \\
3.25 & {1 \over 13} \\
4.75 & {3 \over 52} \\
6.5 & {1 \over 52} \\
 
\end{array}"

"E(X)=(-2)(\dfrac{9}{13})+(-0.75)(\dfrac{1}{13})+(1.25)(\dfrac{1}{13})"

"+(3.25)(\dfrac{1}{13})+(4.75)(\dfrac{3}{52})+(6.5)(\dfrac{1}{52})"

"=-\dfrac{145}{208}"